The 14C activity of some ancient Peruvian corn was found to be 10 disintegrations per minute per gram of C. If present-day plant life shows 15 dpm/g, how old is the Peruvian corn? The halflife of 14C is 5730 years.


1
Expert's answer
2020-07-01T15:43:00-0400
"A=A_0\\cdot 2^{\\frac{t}{T}}\n\\\\\n\n\n\u200b\n\n 15=10\\cdot 2^{\\frac{t}{5730}}\n\n\n\n\u200b\\\\\n\n t=3350\\ years"